OpenStudy (anonymous):

Find all real solutions to the equation f(x)=x^2-5 a) f(x)=4 b) f(x)=-6

OpenStudy (anonymous):

4=x^2-5 9=x^2 x= plus or minus 3 -6=x^2-5 -1=x^2 No real solutions